# A Microscopic Perspective on Moiré Materials

### v1

@@ VIII Outlook
Going forward, we believe that there are three research focuses that are particularly important to the near-term goals of this community. First, our community will benefit from addressing several technical limitations of existing local probes that impede our ability to better understand moiré materials. This includes a focus on new techniques for preserving air-sensitive moiré materials (ex. tMoTe ) in ways compatible with surface-sensitive local probes (ex. STM, KPFM), and new methods for conducting local electromagnetic measurements (ex. nano-SQUID, MIM) at millikelvin temperatures required for stabilizing fragile correlated phases.
@@
Second, our community will benefit from further developing new local probe techniques introduced in the moiré material era. Atomic-scale charge-sensing experiments using the STM, as used to image generalized Wigner crystal states in WSe / WS [105], have great potential in future studies of moiré materials, and might even be capable of quantitative measurements of related physical observables (ex. electronic compressibility) at the nanometer scale. A new nano-SQUID technique can image thermodynamic quantum oscillations in 2D devices to reconstruct local band structure features, which has the potential to provide sensitive measurements of subtle moiré flat band features [203, 204]. The quantum twisting microscope, a revolutionary new measurement tool capable of local 2D wavefunction interference experiments using van der Waals device-structured sensors, has opened up new opportunities to manipulate moiré materials on demand [205]. Further developing this tool’s momentum-space imaging capabilities, and exploring new applications of its novel sensor geometry will support an important new microscopic perspective on moiré materials.
@@
Finally, our community will benefit from continuing to search in moiré materials for exotic quantum phases unrealized elsewhere in Nature, leveraging the many advantages of local probe techniques. Nearly all electronic phases observed in moiré materials to-date have been studied to some degree in other material classes, albeit moiré materials offer highly tunable platforms that enable new investigative approaches. However, we emphasize the importance of finding and understanding novel phases without other material realizations, like the unique time-reversal breaking superconducting state found in twisted cuprate interfaces [206] and the non-equilibrium ferromagnetic and bosonic correlated insulating states found in WSe / WS [207, 208]. Such efforts will support the greatest impact of this budding research field. An exciting prospect is the recent identification of fractional Chern insulators (FCIs) stabilized at zero magnetic field. FCIs have now been reported in tMoTe [22, 23, 70, 71] and in rhombohedral pentalayer graphene aligned to hBN [153], and very recent theoretical studies have suggested that such phases may be stabilized in other moiré materials [209, 210, 211, 212, 213, 214].
@@
Understanding a series of open questions about the nature of FCI states will be a major research focus of local probe experiments in the coming years. To what extent do the recently observed Jain-type FCIs (i.e. , FCIs) parallel Jain states observed in the fractional quantum Hall regime? What kinds of zero-field fractional phases can we realize that are fundamentally different from those realized at high magnetic fields (ex. time-reversal symmetric fractional states, new non-abelian states) [215, 216, 217, 218]? How can we understand FCI states “beyond the Standard Model”, which have no clear fractional quantum Hall analogues (ex. FCI emanating from in MATBG / hBN)? Key to answering these questions will be the microscopic perspective of local probe techniques, which access distinguishable physical observables of these complex ground states and can surpass the limitations of sample inhomogeneity to identify fragile FCI phases in the cleanest, nanometer-scale domains of moiré material devices. Furthermore, certain local probe techniques compatible with top-gated device geometries, including scanning nano-SQUID and atomic-scale charge-sensing using the STM, are distinctly capable of accessing the complete phase diagrams of tMoTe and rhombohedral graphene aligned to hBN, which we expect will provide early microscopic insights into these FCIs.

### v2

@@ VIII Outlook
Going forward, we believe that there are three research focuses that are particularly important to the near-term goals of this community. First, our community will benefit from addressing several technical limitations of existing local probes that impede our ability to better understand moiré materials. This includes a focus on new techniques for preserving air-sensitive moiré materials (ex. tMoTe2 ) in ways compatible with surface-sensitive local probes (ex. STM, KPFM), and new methods for conducting local electromagnetic measurements (ex. nano-SQUID, MIM) at millikelvin temperatures required for stabilizing fragile correlated phases.
@@
Second, our community will benefit from further developing new local probe techniques introduced in the moiré material era. Atomic-scale charge-sensing experiments using the STM, as used to image generalized Wigner crystal states in WSe2 / WS2 [105], have great potential in future studies of moiré materials, and might even be capable of quantitative measurements of related physical observables (ex. electronic compressibility) at the nanometer scale. A new nano-SQUID technique can image thermodynamic quantum oscillations in 2D devices to reconstruct local band structure features, which has the potential to provide sensitive measurements of subtle moiré flat band features [203, 204]. The quantum twisting microscope, a revolutionary new measurement tool capable of local 2D wavefunction interference experiments using van der Waals device-structured sensors, has opened up new opportunities to manipulate moiré materials on demand [205]. Further developing this tool’s momentum-space imaging capabilities, and exploring new applications of its novel sensor geometry will support an important new microscopic perspective on moiré materials.
@@
Finally, our community will benefit from continuing to search in moiré materials for exotic quantum phases unrealized elsewhere in Nature, leveraging the many advantages of local probe techniques. Nearly all electronic phases observed in moiré materials to-date have been studied to some degree in other material classes, albeit moiré materials offer highly tunable platforms that enable new investigative approaches. However, we emphasize the importance of finding and understanding novel phases without other material realizations, like the unique time-reversal breaking superconducting state found in twisted cuprate interfaces [206] and the non-equilibrium ferromagnetic and bosonic correlated insulating states found in WSe2 / WS2 [207, 208]. Such efforts will support the greatest impact of this budding research field. An exciting prospect is the recent identification of fractional Chern insulators (FCIs) stabilized at zero magnetic field. FCIs have now been reported in tMoTe2 [22, 23, 70, 71] and in rhombohedral pentalayer graphene aligned to hBN [153], and very recent theoretical studies have suggested that such phases may be stabilized in other moiré materials [209, 210, 211, 212, 213, 214, 215, 216].
@@
Understanding a series of open questions about the nature of FCI states will be a major research focus of local probe experiments in the coming years. To what extent do the recently observed Jain-type FCIs (i.e. , FCIs) parallel Jain states observed in the fractional quantum Hall regime? What kinds of zero-field fractional phases can we realize that are fundamentally different from those realized at high magnetic fields (ex. time-reversal symmetric fractional states, new non-abelian states) [217, 218, 219, 220]? How can we understand FCI states “beyond the Standard Model”, which have no clear fractional quantum Hall analogues (ex. FCI emanating from in MATBG / hBN)? Key to answering these questions will be the microscopic perspective of local probe techniques, which access distinguishable physical observables of these complex ground states and can surpass the limitations of sample inhomogeneity to identify fragile FCI phases in the cleanest, nanometer-scale domains of moiré material devices. Furthermore, certain local probe techniques compatible with top-gated device geometries, including scanning nano-SQUID and atomic-scale charge-sensing using the STM, are distinctly capable of accessing the complete phase diagrams of tMoTe2 and rhombohedral graphene aligned to hBN, which we expect will provide early microscopic insights into these FCIs.